Q) Consider the following ungrouped data:

41    46     7    46    32     5    14    28    48    49     8    49    48    25    41     8    22    46    40    48

Find the following:

a)    Arithmetic mean

b)   Geometric mean

c)    Harmonic mean

d)   Median

e)    Mode

f)     Range

g)   Mean deviation

h)   Variance

i)     Standard Deviation


1
Expert's answer
2021-05-25T10:12:23-0400
"5, 7, 8, 8, 14, 22, 25, 28, 32, 40,"

"41, 41, 46, 46, 46, 48, 48, 48, 49,49"


a)

Arithmetic mean"=\dfrac{\displaystyle\sum_{i=1}^{20}x_i}{20}=\dfrac{651}{20}=32.55"


b)

Geometric mean"=\sqrt[20]{\displaystyle\prod_{i=1}^{20}x_i}\approx26.3908"


c)

Harmonic mean"=\dfrac{20}{\displaystyle\sum_{i=1}^{20}\dfrac{1}{x_i}}\approx18.8854"

d)

Median = "\dfrac{40+41}{2}=40.5"


e)

Modes are 46 and 48. Each appeared 3 times


f) Range "=49-5=44"


g)  Mean deviation

Mean absolute deviation "MAD"

"MAD=\dfrac{1}{20}\displaystyle\sum_{i=1}^{20}|x_i-\bar{x}|=14.395"


h)  Variance

The variance of a sample is:


"s^2=\dfrac{1}{20-1}\displaystyle\sum_{i=1}^{20}(x_i-\bar{x})^2\approx 270.99736842"


i)  Standard Deviation


"s=\sqrt{s^2}\approx\sqrt{270.99736842}\approx16.4620"